# Heavy Traffic Diffusion Limit for a Closed Queueing Network with Single-Server and Infinite-Server Stations

### v2

@@ 4 Auxiliary Results
This section establishes the existence of (suitably defined) continuous functions that will aid in the proof of Theorem 1 via a continuous mapping argument; see Appendix B for the proofs of the results in this section. To that end, let $\xi\in D^{J}$ and $\zeta\in D^{K}$ be functions such that
@@
Below is a proof of Proposition 1.
Proof of Proposition 1.
It follows that $u_{j}=\mu_{j}^{-1}\psi(z_{j})$ and $x_{j}=\phi(z_{j})$. But then, by the definition of $(\psi,\phi)$ in (1)–(2), it follows that
[…]
Therefore, (57) holds.
We now prove uniqueness. Suppose there are two functions $(x,u,y),(\tilde{x},\tilde{u},\tilde{y})\in D^{2J+K}$ that both satisfy (53)–(57). Then, by (53), for $j\in[J]$ we have that
$\!\!\!\!\!\!\!\!\!\!\!\!\!x_{j}\,=\,z_{j}\,+\,\mu_{j}u_{j}\,\geq\,0\quad\text{and}\quad\tilde{x}_{j}\,=\,\tilde{z}_{j}\,+\,\mu_{j}\tilde{u}_{j}\,\geq\,0,$ (62)
[…]
It then follows from (54), (61), (63), and Lemma 1 that $y_{k}=\tilde{y}_{k}$ for all $k\in[K]$. By uniqueness of $y$, it follows from (61) and (63) that $u_{j}=\tilde{u}_{j}$ for all $j\in[J]$. Finally, by uniqueness of $y$ and $u$, it follows from (61)–(62) that $x_{j}=\tilde{x}_{j}$ for all $j\in[J]$. This completes the proof.∎
@@
Letting $f:=(f_{1},f_{2},f_{3})$, it follows that whenever $(\xi,\zeta)\in D^{J+K}$ satisfies (51)–(52), $f(\xi,\zeta)\in D^{2J+K}$ satisfies (53)–(57). The next result establishes continuity of $f$.
@@ 5.1 Convergence of Fluid Scaled Processes
To prove that $(\bar{\xi}^{n},\bar{\zeta}^{n})\Rightarrow\mathbf{0}$ as $n\rightarrow\infty$, it suffices to show that $\bar{\xi}_{j}^{n}\Rightarrow 0$ as $n\rightarrow\infty$ for all $j\in[J]$ and $\bar{\xi}_{k}^{n}\Rightarrow 0$ as $n\rightarrow\infty$ for all $k\in[K]$; see, e.g., Whitt [26, Theorem 11.4.5]. In turn, it suffices to show that for all $T>0$,
@@
for all $j\in[J]$ and $k\in[K]$; see Lemma 6 in Appendix C for a proof of this claim. By (37)–(38), the triangle inequality, and the fact that $\int_{0}^{t}\bar{\bar{V}}_{k}^{n}(s)\,ds\leq t$ and $T_{j}^{n}(t)\leq t$ for all $t\geq 0$, it follows that for all $T>0$,
@@
By Donsker’s theorem, the functional central limit theorem for renewal processes, the random time change theorem, and the continuous mapping theorem, it is straightforward to show that the right-hand sides of (67) and (68) converge weakly to a nondegenerate limit; see, e.g., Billingsley [8] and Glynn [16]. By this and the fact that $\bar{\xi}_{j}^{n}=n^{-1/2}\hat{\xi}_{j}^{n}$ and $\bar{\zeta}_{k}^{n}=n^{-1/2}\hat{\zeta}_{k}^{n}$, we obtain (66). This is a standard argument, so the detailed proof is omitted.∎
@@
Since $(\bar{\xi}^{n},\bar{\zeta}^{n})\in D^{J+K}$ pathwise, it follows from (39)–(40), (69)–(71), and Proposition 1 that $(\bar{Q}^{n},I^{n},\bar{V}^{n})=(f_{1}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{2}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{3}(\bar{\xi}^{n},\bar{\zeta}^{n}))$. Then, by Proposition 2, Lemma 2, and the continuous mapping theorem, it follows that
@@
To complete the proof, it suffices to show that $\bar{V}\coloneqq f_{3}(\mathbf{0})=\mathbf{0}$, for then it follows from (64)–(65) that $f_{1}(\mathbf{0})=\mathbf{0}$ and $f_{2}(\mathbf{0})=\mathbf{0}$. To that end, it follows from (58), the definition of $\bar{V}$, and the triangle inequality that, for any fixed $T>0$, all $t\in[0,T]$, and all $k\in[K]$, we have
@@
By Gronwall’s inequality (see, e.g., Pang et al. [24, Lemma 4.1]) and (72), it follows that $\max_{k\in[K]}\|\bar{V}_{k}\|_{T}=0.$ Since $T$ was arbitrary, it follows that $\bar{V}\equiv 0$. This completes the proof.∎
@@
By definition, $T^{n}=e-I^{n}$; see (13). The result then follows by Lemma 3 since $I^{n}\Rightarrow\mathbf{0}$ as $n\rightarrow\infty$.∎
@@ 5.2 Convergence of Diffusion Scaled Processes
where $B_{kj}$, $\tilde{B}_{jk}$, $B_{k}$, and $\tilde{B}_{j}$ are independent standard Brownian motions. (These convergence arguments are routine and therefore omitted for brevity.) Moreover, the mapping $H:D\rightarrow D$ defined by
$H(x)(t)\coloneqq\!\int_{0}^{t}\!x(s)\,ds,\quad(x,t)\,\in\,D\,\times\,[0,\infty),$
is continuous in the Skorokhod topology (see, e.g., Pang et al. [24, page 229]), which implies that
@@
By the above weak convergence results, Whitt [26, Theorems 11.4.4 and 11.4.5], Assumption 2, and the fact that all stochastic primitives are independent, it follows that the (joint) processes $(\hat{Q}^{n}(0),\,\hat{V}^{n}(0),\,\hat{\Psi}^{n},\,\hat{\Phi}^{n},\,\hat{N}^{n},\dots,\,\hat{M}^{n})$ and $(T^{n},\,\bar{\bar{V}}^{n},\,\bar{N}^{n},\,\bar{M}^{n})$ converge weakly as $n\rightarrow\infty$ to their appropriate limits. From this, (37)–(38), the random time change theorem, and the continuous mapping theorem, it follows that $(\hat{\xi}^{n},\hat{\zeta}^{n})$ converges weakly as $n\rightarrow\infty$ to $(J+K)$-dimensional Brownian motion $(\xi^{*},\zeta^{*})$ with initial state $(Q(0),V(0))$ and covariance matrix $\Sigma$ given by (42)–(46). Because it is straightforward, albeit tedious, to derive the entries of the covariance matrix $\Sigma$, we omit the details. ∎
@@ A.1 Existence of an element in $D_{K}^{T}$ satisfying (58) for all $t\in[0,T]$
We prove existence via the method of successive approximations; see, e.g., Reed and Ward [25] for a similar proof approach. To be more specific, we show that the sequence formed by this method is Cauchy in $D_{K}^{T}$ under the sup norm, and then argue that the limit of the sequence (which exists by completeness of $D_{K}^{T}$) satisfies (58).
@@
Then, for $y,\tilde{y}\in D_{T}^{K}$, it follows by the definition of $L$, the triangle inequality, and (3) that
@@
implying that that $L$ is Lipschitz continuous. It then follows from (78)–(79) that
@@
for all $k\in[K]$, where the first inequality follows from (3) and (78). By performing a similar estimate of $\|y_{l}^{n-1}-y_{l}^{n-2}\|_{\delta}$ in the above display, it follows that
@@ Appendix B Proof of Proposition 2
Suppose that $(\xi^{n},\zeta^{n})\rightarrow(\xi,\zeta)$ in $D^{J+K}$ as $n\rightarrow\infty$ and let $\tilde{T}>0$ be a continuity point of $f_{3}(\xi,\zeta)$. To prove that $f_{3}$ is continuous with respect to the Skorokhod $J_{1}$ topology, we must show that
$\lim_{n\rightarrow\infty}d_{\tilde{T}}^{K}\big(f_{3}(\xi^{n},\zeta^{n})|_{[0,\tilde{T}]},f_{3}(\xi,\zeta)|_{[0,\tilde{T}]}\big)\,=\,0;$
see the technical preliminaries given in Section 1. Since $(\xi,\zeta)\in D^{J+K}$, it has at most countably many points of discontinuity; see, e.g., Ethier and Kurtz [14, Lemma 5.1]. It follows that there exists some $T>\tilde{T}$ that is a continuity point of $(\xi,\zeta)$. Therefore, by Billingsley [8, Lemma 1, page 167], it suffices to show that
@@
where the first inequality holds by (3) and the fact that
@@
see, e.g., Whitt [26, Lemma 13.5.2]. Combining these estimates, it follows from (87)–(89) that
@@
for all $0\leq t\leq T$. By Gronwall’s inequality (see, e.g., Pang et al. [24, Lemma 4.1]) and the above displayed inequality, it follows that
@@
Appendix C Miscellaneous Proofs
@@
see, e.g., Whitt [26, Chapter 3, Section 3]. But observe that

### v3

@@ 4 Auxiliary Results
This section establishes the existence of (suitably defined) continuous functions that will aid in the proof of Theorem 1 via a continuous mapping argument. To that end, let $\xi\in D^{J}$ and $\zeta\in D^{K}$ be functions such that
@@
Let $f:=(f_{1},f_{2},f_{3})$. Then, whenever $(\xi,\zeta)\in D^{J+K}$ satisfies (49)–(50), $f(\xi,\zeta)\in D^{2J+K}$ satisfies (51)–(55). The next result establishes continuity of $f$.
@@ 5.1 Convergence of Fluid Scaled Processes
To prove that $(\bar{\xi}^{n},\bar{\zeta}^{n})\Rightarrow\mathbf{0}$ as $n\rightarrow\infty$, it suffices to show that $\bar{\xi}_{j}^{n}\Rightarrow 0$ as $n\rightarrow\infty$ for all $j\in[J]$ and $\bar{\xi}_{k}^{n}\Rightarrow 0$ as $n\rightarrow\infty$ for all $k\in[K]$; see, e.g., Whitt [25, Theorem 11.4.5]. In turn, it suffices to show that for all $T>0$,
@@
for all $j\in[J]$ and $k\in[K]$; see Lemma 6 in Appendix D for a proof of this claim. By (35)–(36), the triangle inequality, and the fact that $\int_{0}^{t}\bar{\bar{V}}_{k}^{n}(s)\,ds\leq t$ and $T_{j}^{n}(t)\leq t$ for all $t\geq 0$, it follows that for all $T>0$,
@@
By Donsker’s theorem, the functional central limit theorem for renewal processes, the random time change theorem, and the continuous mapping theorem, it is straightforward to show that the right-hand sides of (60) and (61) converge weakly to nondegenerate limits; see, e.g., Billingsley [6] and Glynn [14]. By this and the fact that $\bar{\xi}_{j}^{n}=n^{-1/2}\hat{\xi}_{j}^{n}$ and $\bar{\zeta}_{k}^{n}=n^{-1/2}\hat{\zeta}_{k}^{n}$, we obtain (59). This is a standard argument, so the detailed proof is omitted.∎
@@
Since $(\bar{\xi}^{n},\bar{\zeta}^{n})\in D^{J+K}$ pathwise, it follows from (37)–(38), (62)–(64), and Proposition 1 that $(\bar{Q}^{n},I^{n},\bar{V}^{n})=(f_{1}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{2}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{3}(\bar{\xi}^{n},\bar{\zeta}^{n}))$. Then, by Proposition 2, Lemma 2, and the continuous mapping theorem,
@@
It now suffices to show that $\bar{V}\coloneqq f_{3}(\mathbf{0})=\mathbf{0}$, for then it follows from (57)–(58) that $f_{1}(\mathbf{0})=\mathbf{0}$ and $f_{2}(\mathbf{0})=\mathbf{0}$. To that end, it follows from (56), the definition of $\bar{V}$, and the triangle inequality that, for any fixed $T>0$, all $t\in[0,T]$, and all $k\in[K]$, we have
@@
By Gronwall’s inequality (see, e.g., Pang et al. [23, Lemma 4.1]) and (65), it follows that $\max_{k\in[K]}\|\bar{V}_{k}\|_{T}=0.$ Since $T$ was arbitrary, it follows that $\bar{V}\equiv 0$.∎
@@
By the definition in (11), $T^{n}=e-I^{n}$. The result then follows by Lemma 3 since $I^{n}\Rightarrow\mathbf{0}$ as $n\rightarrow\infty$.∎
@@ 5.2 Convergence of Diffusion Scaled Processes
where $B_{kj}$, $\tilde{B}_{jk}$, $B_{k}$, and $\tilde{B}_{j}$ are independent standard Brownian motions. (These convergence arguments are routine and therefore omitted for brevity.) Furthermore, the mapping $H:D\rightarrow D$, defined by $H(x)(t)\coloneqq\int_{0}^{t}\!x(s)\,ds$ for $(x,t)\in D\times[0,\infty)$, is continuous in the Skorokhod $J_{1}$ topology (see, e.g., Pang et al. [23, page 229]), which implies that
@@
By the above weak convergence results, Whitt [25, Theorems 11.4.4 and 11.4.5], Assumption 2, and the independence of the stochastic model primitives, it follows that the (joint) processes $(\hat{Q}^{n}(0),\,\hat{V}^{n}(0),\,\hat{\Psi}^{n},\,\hat{\Phi}^{n},\,\hat{N}^{n},\dots,\,\hat{M}^{n})$ and $(T^{n},\,\bar{\bar{V}}^{n},\,\bar{N}^{n},\,\bar{M}^{n})$ converge weakly as $n\rightarrow\infty$ to their appropriate limits. From this, (35)–(36), the random time change theorem, and the continuous mapping theorem, it follows that $(\hat{\xi}^{n},\hat{\zeta}^{n})$ converges weakly as $n\rightarrow\infty$ to $(J+K)$-dimensional Brownian motion $(\xi^{*},\zeta^{*})$ with initial state $(Q(0),V(0))$ and covariance matrix $\Sigma$ given by (40)–(44). Because it is straightforward, albeit tedious, to derive the entries of the covariance matrix $\Sigma$, we omit the details. ∎
@@ A.1 Existence of an element in $D_{K}^{T}$ satisfying (56) for all $t\in[0,T]$
We prove existence via the method of successive approximations; see, e.g., Reed and Ward [24] for a similar proof approach. In particular, we construct a sequence that is Cauchy in $D_{K}^{T}$ under the sup norm, and then argue that the limit of the sequence (which exists by completeness of $D_{K}^{T}$) satisfies (56).
@@
Then, for $y,\tilde{y}\in D_{T}^{K}$, it follows by the definition of $L$, the triangle inequality, and Whitt [25, Lemma 13.5.1] that
@@
implying that $L$ is Lipschitz continuous. It then follows from (71)–(72) that
@@
for all $k\in[K]$, where the first inequality follows from Whitt [25, Lemma 13.5.1] and (71). By performing a similar estimate of $\|y_{l}^{n-1}-y_{l}^{n-2}\|_{\delta}$ in the above display, it follows that
@@ Appendix B Proof of Proposition 1
Thus, $u_{j}=\mu_{j}^{-1}\psi(z_{j})$ and $x_{j}=\phi(z_{j})$. But, by (1)–(2), it follows that $\int_{0}^{\infty}\mathds{1}_{\{x_{j}>0\}}\,d(u_{j})(t)=0.$ Therefore, (55) holds.
We now prove uniqueness. Suppose $(x,u,y),(\tilde{x},\tilde{u},\tilde{y})\in D^{2J+K}$ both satisfy (51)–(55). Then, by (51), we have that
[…]
It then follows from (52), (80), (82), and Lemma 1 that $y_{k}=\tilde{y}_{k}$ for all $k\in[K]$. By uniqueness of $y$, it follows from (80) and (82) that $u_{j}=\tilde{u}_{j}$ for all $j\in[J]$. Finally, by uniqueness of $y$ and $u$, it follows from (80)–(81) that $x_{j}=\tilde{x}_{j}$ for all $j\in[J]$. This completes the proof.
Appendix C Proof of Proposition 2
@@
In the Skorokhod $J_{1}$ topology (see, e.g., Billingsley [6] and Whitt [25]), a sequence $\{x^{n}\}_{n=1}^{\infty}$ in $D^{k}$ converges to an element $x\in D^{k}$, written $x^{n}\rightarrow x$, as $n\rightarrow\infty$, if $d_{T}^{k}(x^{n}|_{[0,T]},x|_{[0,T]})\rightarrow 0$ as $n\rightarrow\infty$ for all continuity points $T>0$ of $x$, where $d_{T}^{k}:D_{T}^{k}\,\times\,D_{T}^{k}\rightarrow[0,\infty)$ is given by
where $e:[0,T]\rightarrow[0,T]$ is the identity map, and
$\Lambda_{T}\coloneqq\big\{\lambda$ $:[0,T]\rightarrow[0,T]$
$|\,\lambda\text{ is an increasing homeomorphism}\big\}.$
To that end, suppose that $(\xi^{n},\zeta^{n})\rightarrow(\xi,\zeta)$ in $D^{J+K}$ as $n\rightarrow\infty$ and let $\tilde{T}>0$ be a continuity point of $f_{3}(\xi,\zeta)$. To prove that $f_{3}$ is continuous with respect to the Skorokhod $J_{1}$ topology, we must show that
$\lim_{n\rightarrow\infty}d_{\tilde{T}}^{K}\big(f_{3}(\xi^{n},\zeta^{n})|_{[0,\tilde{T}]},f_{3}(\xi,\zeta)|_{[0,\tilde{T}]}\big)\,=\,0.$
Since $(\xi,\zeta)\in D^{J+K}$, it has at most countably many points of discontinuity; see, e.g., Ethier and Kurtz [12, Lemma 5.1]. It follows that there exists some $T>\tilde{T}$ that is a continuity point of $(\xi,\zeta)$. Therefore, by Billingsley [6, Lemma 1, page 167], it suffices to show that
@@
where the first inequality holds by Whitt [25, Lemma 13.5.1] and the fact that
@@
see, e.g., Whitt [25, Lemma 13.5.2]. Combining these estimates, it follows from (85)–(87) that
@@
for all $0\leq t\leq T$. By Gronwall’s inequality (see, e.g., Pang et al. [23, Lemma 4.1]) and the above displayed inequality, it follows that
@@
Appendix D Miscellaneous Proofs
@@
see, e.g., Whitt [25, Chapter 3, Section 3]. But observe that